Solve the system(Substitution method recommended)
x+y=32
y=x^2-4x+4

Answer by ewatrrr(24785)


Hi
y = 32-x
32-x = x^2-4x+4
x^2 -3x - 28 = 0
(x-7)(x+4) = 0, x = -4, 7 and y = 36, 25
(-4, 36) and (7,25) are the solutions for this System